# GlaLSTM: A Concurrent LSTM Stream Framework for Glaucoma Detection via Biomarkers Mining

### v2

@@ Abstract
Glaucoma is a complex group of eye diseases marked by optic nerve damage, commonly linked to elevated intraocular pressure and biomarkers like retinal nerve fiber layer thickness. Understanding how these biomarkers interact is crucial for unraveling glaucoma’s underlying mechanisms. In this paper, we propose GlaLSTM, a novel concurrent LSTM stream framework for glaucoma detection, leveraging latent biomarker relationships. Unlike traditional CNN-based models that primarily detect glaucoma from images, GlaLSTM provides deeper interpretability, revealing the key contributing factors and enhancing model transparency. This approach not only improves detection accuracy but also empowers clinicians with actionable insights, facilitating more informed decision-making. Experimental evaluations confirm that GlaLSTM surpasses existing state-of-the-art methods, demonstrating its potential for both advanced biomarker analysis and reliable glaucoma detection.
Keywords Glaucoma Biomarker  $\cdot$ Data Mining  $\cdot$ Long short-term memory  $\cdot$ Latent Relationship Exploration  $\cdot$ Glaucoma Classification
1 INTRODUCTION
@@
Recently, artificial intelligence (AI) and machine learning (ML) techniques have been increasingly integrated into glaucoma research to enhance early detection [1, 5, 4, 6]. Given that delayed medical attention and treatment can lead to irreversible blindness, researchers have focused on developing AI-driven solutions to improve diagnostic accuracy [7, 8, 9]. Most of these cross-disciplinary efforts are rooted in the domain of computer vision, particularly in image detection and segmentation techniques [1, 5, 4]. Meanwhile, convolutional neural networks (CNNs) [10, 4, 11, 12, 6] and attention mechanisms [13, 14] have demonstrated impressive performance. However, despite their accuracy, most ML-based models function as "black boxes," lacking explainability. This opacity raises concerns among clinicians, vision scientists, and patients, who may hesitate to fully trust automated algorithms with unclear decision-making processes.
@@ 3 Methodology
As illustrated in Fig. 1, the concurrent LSTM stream framework comprises two LSTMs and integrates co-attention mechanism [13, 14], where one LSTM and co-attention are employed twice within the structure. While decoding the first co-attention aims to capture features of original input biomakrer and its output is ($T_{t}^{1}$), sequence and image but the second one focuses on random order input and its output is ($T_{t}^{2}$). LSTM is a time-series model where earlier inputs gradually lose influence due to the forgetting mechanism, affecting long-term dependencies. We designed a three-stage LSTM framework. Two LSTMs process the same data in different orders to ensure uniform biomarker priority. By comparing their outputs, we identify common and distinct biomarkers. A third LSTM then integrates both sequences, doubling common biomarkers while keeping distinct ones single for better feature extraction.
@@
4 Data and Evaluation Matrix
4.1 Data
@@
Table 1: One Sample from UTSW
Category Biomarker Details
@@
For private data from UTSW, as shown in Table. 1, this is a sample collected from a healthy person. For its preprocessing, there are three types of biomarkers themselves: Within Normal, Borderline and Outside Normal. For instance, A-G is the borderline and its label is "Borderline" or 0.5, S-G is outside normal, labeled "False" or 0, and AR is nomal, labeled "True" or 1. Among them, Borderline is a critical situation, which is normal in medical terms, but its value is close to the pathology. Unless explicitly stated as normal, it is classified into a third category. In addition, we also considered some other non-biomarker influencing factors: age, gender, race, region, etc. Finally, we have two glaucoma experts from UTSW in our team who can help us with annotations.
Note: retinal nerve fiber layer = RNFL (Average RNFL = AR; Superior RNFL = SR; Inferior RNFL = IR; Intra Eye (S-I) = I-ER); optic nerve head = ONH; Cup/Disc Area Ratio = A-O; Cup/Disc V. Ratio = V-O; Cup/Disc H. Ratio = H-O; Rim Area = RA; Disc Area = DA; Cup Volume = CVO; ganglion cell complex = GCC; Average GCC = A-G; Superior GCC = S-G; Inferior FCC = I-F; Intra Eye (S-I) = I-EG (S-I); right eye = OD and left eye = OS. IE (OD-OS) is a condition that affects both the right and left eyes. Except for GCC, RNFL and ONH, the remaining abbreviations are only used in this paper. The number of training and test samples is 500 and 100 respectively. IOP is an additional input.
For publuc data, as shown in Table. 2, compared with data in the Table. 1, the amount of data has been reduced significantly. For its preprocessing to features like optic_disc_size, cup_to_disc_ratio and neuroretinal_rim which have specific value or description, we will refer to the judgment criteria for glaucoma and give "True" or "False".
Table 2: One Sample from Public Dataset
@@
Note: "neuroretinal_rim" has been simplified and the original version is "thinned in all regions, especially inferior and superior." "glaucoma_risk_assessment" and "confidence_level" are not biomarkers. They are pre-judged by humans to predict the probability of glaucoma and their own confidence in that judgment. The annotation of this sample is "glaucoma". The number of training and test samples is 589 and 100 respectively.
Unlike Table. 1, glaucoma_risk_assessment and confidence_level are indicators that are prejudged, which will greatly affect the potential relationship mining of various biomarkers. Moreover, the OS and OD parameters, representing the left and right eyes, along with the IE (OD-OS) parameter in the data from Table. 1, are three times the number of parameter names listed, whose variable value (32*3, single eye; 48*3, double eyes) is 6 or 4 times greater than those (12*2) in Table. 2 (in this case, "Borderline" is reguarded as nomal, or the third category, it is 4 or 2.7 times). In this study, the experiment we designed for UTSW is for both eyes.
4.2 Evaluation Matrix
To evaluate the directional accuracy of our detection, we frame this assessment as a binary classification task. This approach allows us to measure the model’s ability to correctly identify market trends using a confusion matrix, which provides a comprehensive performance evaluation through the following metrics:
  • True Positive (TP): the TP indicates that Class A is correctly identified as belonging to Class A.
  • True Negative (TN): this matrix indicates that Class A is correctly identified as not belonging to Class A.
  • False Positive (FP): it indicates that Class A is not correctly identified as belonging to Class A.
  • False Negative (FN): it indicates that Class A is not correctly identified as not belonging to Class A.
Especially in classification tasks, three key evaluation metrics are precision, accuracy, and recall. For precision, it measures whether results are actually positive. High precision means fewer FP. Its formula is shown below in Eq. 6:
$\begin{split}\text{Precision}=\frac{TP}{TP+FP}\end{split}$ (6)
For accuracy, it measures the overall correctness of the model by calculating the percentage of correctly classified samples (both positive and negative). Its formula is shown below in Eq. 7:
$\begin{split}\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}\end{split}$ (7)
For recall, it measures how many actual positive cases were correctly identified by the model. Its formula is shown below in Eq. 8:
$\begin{split}\text{Recall}=\frac{TP}{TP+FN}\end{split}$ (8)
5 Experimental Result and Analysis
5.1 Classification Results
As shown in Table. 3, it presents the experimental results of various models evaluated on public data, comparing their performance in terms of Precision, Recall, and F1-score, both with and without the "Prejudge" mechanism. Enabling Prejudge consistently improves all models’ performance. For instance, RNN-LSTM’s F1-score rises from 0.694 without Prejudge to 0.824 with full Prejudge, and similar improvements are observed for Random Forest (0.717 to 0.831) and Bi-LSTM (0.715 to 0.827). Notably, Decision Tree and XGBoost also show substantial gains in F1-score with full Prejudge, reaching 0.811 and 0.879, respectively. Among the models, XGBoost demonstrates more stable performance metrics compared to GPT-4o, although the overall results of the two are roughly comparable. Our proposed model achieves the highest overall performance, with an F1-score of 0.913 under full Prejudge, highlighting its superior ability to handle classification tasks.
Table 3: Experimental Results on public data.
@@ 5.1 Classification Results
Note: if glaucoma_risk_assessment and confidence_level are enabled, giving them a higher weight, the thief is ✓; otherwise, the weight is the same as other weights, then $\sqrt{}\mkern-9.0mu{\smallsetminus}$; if not enabled, then $\times$.
Table 4: Experimental Results on data from UTSW
@@
As shown in Table 4, the concurrent LSTM stream framework achieves the highest F1-score (0.857) with all conditions enabled, demonstrating its robustness. XGBoost (0.819) and GPT-4o (0.827) also show strong, stable performance. In contrast, RNN-LSTM, Random Forest, and Bi-LSTM perform better with fewer conditions but decline under full settings, indicating sensitivity to complexity and noise. Decision Tree performs consistently, peaking at 0.795 with all conditions. These results highlight the proposed model’s effectiveness and the stability of XGBoost and GPT-4o in diverse settings.
Note: "Random" means that ignoring the order of IOP, all biomarkers were initially randomized ($\times$) or on the contrary, IOP presets the highest priority (✓). "Borderline" means that biomarkers with label ’Borderline’ are uniformly classified as the third category (✓), otherwise they are classified as the positive category ($\times$). "Others" means that if some of the non-biological marker information mentioned earlier, such as age, gender, etc., is taken into account (✓), otherwise not ($\times$).
@@
Figure 2: The following statistics pertain to the five most influential biomarkers (UTSW), displayed horizontally in descending order of influence from left to right. Amony them, the symbols below each image are for the three categories in Table. 4: "Random", "Borderline" and "Others".
5.2 Hidden Relationship Mining
@@
Figure 3: The following statistics pertain to the five most influential biomarkers (public data), displayed horizontally in descending order of influence from left to right. Amony them, the symbol below each image are for the category in Table. 3: "Prejudge".
As shown in Fig. 3, when "Prejudge" is marked with ✓, both glaucoma_risk_assessment and confidence_level are included. However, as pre-added labels rather than biomarkers, they are excluded from ranking. We remove them accordingly. In secondary experiments, neuroretinal_rim consistently ranked among the top features. This optic disc structure is crucial for assessing optic nerve health, particularly in glaucoma. Similarly, optic_disc_size, which varies among individuals, is a key parameter in optic nerve evaluation. The prominence of these biomarkers demonstrates our method’s effectiveness in identifying the most influential features for classification.
5.3 Analysis
@@
Our analysis reveals that as the number of calculated variables increases, the model must balance more factors, resulting in lower accuracy when the number of training iterations remains constant. In contrast, if the model is deliberately given a training bias by assigning a higher initial weight to "Prejudge (✓)," the accuracy of the models improves significantly, as shown in Table. 3. These results demonstrate the significant impact of Prejudge on improving model accuracy and emphasize the effectiveness of the proposed framework. Additionally, we observed that when the "Prejudge" condition is assigned the same weight as other biomarkers, the performance of all models declines in the experimental results.
@@
Table 4 Analysis: Enabling the Borderline indicator led to a significant performance drop for RNN-LSTM, Random Forest, and Bi-LSTM, while improving other models. This decline stems from overfitting tendencies, Random Forest struggles with extrapolation, and LSTMs require more data to retain distant sample features. In contrast, decision trees excel due to intrinsic feature selection and multi-type data handling but are highly sensitive to data fluctuations, causing result instability. XGBoost mitigates this via iterative boosting, regularization, and automatic tree pruning, enhancing stability and performance. Unlike decision trees, XGBoost also optimally handles missing values, making it a more robust choice.
@@
Overall, compared to Table 3, the values in Table 4 are generally lower. The key factor is Borderline, when set to "$\times$", the magnitude increases only slightly per case, regardless of being "True" or "False." However, when set to ✓, it effectively introduces a new major category, neutral samples. Treating the process as a multivariate function, this addition corresponds to a new dimension, significantly increasing computational complexity by an entire order of magnitude. Concurrent LSTM stream framework excelled across all metrics, especially after data shuffling, enhancing generalization and accuracy. Using two LSTMs as head inputs broadens feature learning, preventing local optima. Different input orders improve sampling, reducing overfitting. Compared to GPT-4o’s Prompt Project results, our approach is more stable and noise-resistant.
@@
Also, several details are noteworthy for classification: Fig. 2(h) shows that ethnicity under the "Others" category appears once, reflecting glaucoma’s racial differences, consistent with existing literature [2, 17, 3]. Additionally, when "Borderline" changes to $\times$, the model integrates more variables, balancing IOP’s significance by introducing additional factors. While elevated IOP is a primary glaucoma marker, other elements—such as genetics, lifestyle, and regional history—also play crucial roles. Thus, despite IOP’s high frequency in the experiment, the underlying contributing factors warrant deeper consideration.
6 Conclusion
@@
Acknowledgments
This research work is partially sponsored by NASA 80NSSC22K0144, UTSW GMO No.241213, and NIH 1R01AG083179-01.

### v3

# GlaLSTM: A Concurrent LSTM Stream Framework for Glaucoma Detection via Biomarker Relationship Mining

@@ Abstract
Glaucoma is a leading cause of irreversible blindness globally, commonly linked to elevated intraocular pressure and biomarkers like retinal nerve fiber layer thickness. Understanding how these biomarkers interact is crucial for unraveling glaucoma’s underlying mechanisms. In this paper, we propose GlaLSTM, a novel concurrent LSTM stream framework for glaucoma detection, leveraging latent biomarker relationships. Unlike traditional CNN-based models that primarily detect glaucoma from images, GlaLSTM provides deeper interpretability, revealing the key contributing factors and enhancing model transparency. This approach not only improves detection accuracy but also empowers clinicians with actionable insights, facilitating more informed decision-making. Experimental evaluations confirm that GlaLSTM surpasses existing state-of-the-art methods, demonstrating its potential for both advanced biomarker analysis and reliable glaucoma detection.
Keywords— Glaucoma Biomarker, Data Mining, Long short-term memory, Latent Relationship Exploration and Classification
I INTRODUCTION
@@
Recently, artificial intelligence (AI) and machine learning (ML) techniques have been increasingly integrated into glaucoma research to enhance early detection [1, 4, 5, 27]. Given that delayed medical attention and treatment can lead to irreversible blindness, researchers have focused on developing AI-driven solutions to improve diagnostic accuracy [17, 28, 31]. Most of these cross-disciplinary efforts are rooted in the domain of computer vision, particularly in image detection and segmentation techniques [1, 4, 5]. Meanwhile, convolutional neural networks (CNNs) [9, 5, 8, 10, 27] and attention mechanisms [6, 32] have demonstrated impressive performance. However, despite their accuracy, most ML-based models function as “black boxes,” lacking explainability. This opacity raises concerns among clinicians, vision scientists, and patients, who may hesitate to fully trust automated algorithms with unclear decision-making processes.
@@ III Methodology
As illustrated in Fig. 1, the concurrent LSTM stream framework comprises two LSTMs and integrates co-attention mechanism [6, 32], where one LSTM and co-attention are employed twice within the structure.
While decoding the first co-attention aims to capture features of original input biomarker and its output is $T_{t}^{1}$, sequence and image but the second one focuses on random order input and its output is $T_{t}^{2}$.
LSTM is a time-series model where earlier inputs gradually lose influence due to the forgetting mechanism, affecting long-term dependencies. We designed a three-stage LSTM framework. Two LSTMs process the same data in different orders to ensure uniform biomarker priority. By comparing their outputs, we identify common and distinct biomarkers. A third LSTM then integrates both sequences, doubling common biomarkers while keeping distinct ones single for better feature extraction.
@@
IV Data and its Pre-processing
@@
TABLE I: One Sample from UTSW
Category Biomarker Detail
@@
  • 1retinal nerve fiber layer = RNFL (Average RNFL = AR; Superior RNFL = SR; Inferior RNFL = IR; Intra Eye (S-I) = I-ER)
  • 2optic nerve head = ONH (Cup/Disc Area Ratio = A-O; Cup/Disc V. Ratio = V-O; Cup/Disc H. Ratio = H-O; Rim Area = RA; Disc Area = DA; Cup Volume = CVO)
  • 3ganglion cell complex = GCC (Average GCC = A-G; Superior GCC = S-G; Inferior FCC = I-F; Intra Eye (S-I) = I-EG (S-I))
  • 4IE (OD-OS) is a condition that affects both the right and left eyes.
  • 5right eye = OD and left eye = OS
  • 6Except for GCC, RNFL and ONH, the remaining abbreviations are only used in this paper.
  • 7The number of training and test samples is 500 and 100 respectively.
  • 8IOP is an additional input.
For private data from UTSW, as shown in Table. I, this is a sample collected from a healthy person. For its preprocessing, there are three types of biomarkers themselves: Within Normal, Borderline and Outside Normal. For instance, A-G is the borderline and its label is “Borderline” or 0.5, S-G is outside normal, labeled “False” or 0, and AR is nomal, labeled “True” or 1. Among them, Borderline is a critical situation, which is normal in medical terms, but its value is close to the pathology. Unless explicitly stated as normal, it is classified into a third category. In addition, we also considered some other non-biomarker influencing factors: age, gender, race, region, etc. Finally, we have two glaucoma experts from UTSW in our team who can help us with annotations.
TABLE II: One Sample from Public Dataset
@@
  • 1“neuroretinal_rim” has been simplified and the original version is thinned in all regions, especially inferior and superior.”
  • 2“glaucoma_risk_assessment” and“confidence_level” are not biomarkers. They are pre-judged by humans to predict the probability of glaucoma and their own confidence in that judgment.
  • 3The annotation of this sample is “glaucoma”.
  • 4The number of training and test samples is 589 and 100 respectively.
For public data, as shown in Table. II, compared with data in the Table. I, the amount of data has been reduced significantly. For its preprocessing to features like optic_disc_size, cup_to_disc_ratio and neuroretinal_rim which have specific value or description, we will refer to the judgment criteria for glaucoma and give “True” or “False”. Unlike Table. I, glaucoma_risk_assessment and confidence_level are indicators that are prejudged, which will greatly affect the potential relationship mining of various biomarkers. Moreover, the OS and OD parameters, representing the left and right eyes, along with the IE (OD-OS) parameter in the data from Table. I, are three times the number of parameter names listed, whose variable value (32*3, single eye; 48*3, double eyes) is 6 or 4 times greater than those (12*2) in Table. II (in this case, “Borderline” is reguarded as nomal, or the third category, it is 4 or 2.7 times). In this study, the experiment we designed for UTSW is for both eyes.
V Experimental Result and Analysis
V-A Classification Results
TABLE III: Experimental Results on Public Data
@@ V-A Classification Results
  • 1if glaucoma_risk_assessment and confidence_level are enabled, giving them a higher weight, the thief is ✓; otherwise, the weight is the same as other weights, then $\sqrt{}\mkern-9.0mu{\smallsetminus}$; if not enabled, then $\times$.
As shown in Table. III, it presents the experimental results of various models evaluated on public data, comparing their performance in terms of Precision, Recall, and F1-score, both with and without the “Prejudge” mechanism. Enabling Prejudge consistently improves all models’ performance. For instance, RNN-LSTM’s F1-score rises from 0.694 without Prejudge to 0.824 with full Prejudge, and similar improvements are observed for Random Forest (0.717 to 0.831) and Bi-LSTM (0.715 to 0.827). Notably, Decision Tree and XGBoost also show substantial gains in F1-score with full Prejudge, reaching 0.811 and 0.879, respectively. Among the models, XGBoost demonstrates more stable performance metrics compared to GPT-4o, although the overall results of the two are roughly comparable. Our proposed model achieves the highest overall performance, with an F1-score of 0.913 under full Prejudge, highlighting its superior ability to handle classification tasks.
TABLE IV: Experimental Results on Data from UTSW
@@
  • 1“Random” means that ignoring the order of IOP, all biomarkers were initially randomized ($\times$) or on the contrary, IOP presets the highest priority (✓).
  • 2“Borderline” means that biomarkers with label ‘Borderline’ are uniformly classified as the third category (✓), otherwise they are classified as the positive category ($\times$).
  • 3“Others” means that if some of the non-biological marker information mentioned earlier, such as age, gender, etc., is taken into account (✓), otherwise not ($\times$).
As shown in Table IV, the concurrent LSTM stream framework achieves the highest F1-score (0.857) with all conditions enabled, demonstrating its robustness. XGBoost (0.819) and GPT-4o (0.827) also show strong, stable performance. In contrast, RNN-LSTM, Random Forest, and Bi-LSTM perform better with fewer conditions but decline under full settings, indicating sensitivity to complexity and noise. Decision Tree performs consistently, peaking at 0.795 with all conditions. These results highlight the proposed model’s effectiveness and the stability of XGBoost and GPT-4o in diverse settings.
@@
Figure 2: The following statistics pertain to the five most influential biomarkers (UTSW), displayed horizontally in descending order of influence from left to right. Amony them, the symbols below each image are for the three categories in Table. IV: “Random”, “Borderline” and “Others”.
V-B Hidden Relationship Mining
@@
Figure 3: The following statistics pertain to the five most influential biomarkers (public data), displayed horizontally in descending order of influence from left to right. Amony them, the symbol below each image are for the category in Table. III: ”Prejudge”.
V-C Analysis
@@
Our analysis reveals that as the number of calculated variables increases, the model must balance more factors, resulting in lower accuracy when the number of training iterations remains constant. In contrast, if the model is deliberately given a training bias by assigning a higher initial weight to “Prejudge (✓),” the accuracy of the models improves significantly, as shown in Table. III. These results demonstrate the significant impact of Prejudge on improving model accuracy and emphasize the effectiveness of the proposed framework. Additionally, we observed that when the “Prejudge” condition is assigned the same weight as other biomarkers, the performance of all models declines in the experimental results.
@@
Table IV Analysis: Enabling the Borderline indicator led to a significant performance drop for RNN-LSTM, Random Forest, and Bi-LSTM, while improving other models. This decline stems from overfitting tendencies, Random Forest struggles with extrapolation, and LSTMs require more data to retain distant sample features. In contrast, decision trees excel due to intrinsic feature selection and multi-type data handling but are highly sensitive to data fluctuations, causing result instability. XGBoost mitigates this via iterative boosting, regularization, and automatic tree pruning, enhancing stability and performance. Unlike decision trees, XGBoost also optimally handles missing values, making it a more robust choice.
@@
Overall, compared to Table III, the values in Table IV are generally lower. The key factor is Borderline, when set to $\times$”, the magnitude increases only slightly per case, regardless of being “True” or “False.” However, when set to ✓, it effectively introduces a new major category, neutral samples. Treating the process as a multivariate function, this addition corresponds to a new dimension, significantly increasing computational complexity by an entire order of magnitude. Concurrent LSTM stream framework excelled across all metrics, especially after data shuffling, enhancing generalization and accuracy. Using two LSTMs as head inputs broadens feature learning, preventing local optima. Different input orders improve sampling, reducing overfitting. Compared to GPT-4o’s Prompt Project results, our approach is more stable and noise-resistant.
@@
Also, several details are noteworthy for classification: Fig. 2(h) shows that ethnicity under the ”Others” category appears once, reflecting glaucoma’s racial differences, consistent with existing literature [14, 16, 18]. Additionally, when “Borderline” changes to $\times$, the model integrates more variables, balancing IOP’s significance by introducing additional factors. While elevated IOP is a primary glaucoma marker, other elements—such as genetics, lifestyle, and regional history—also play crucial roles. Thus, despite IOP’s high frequency in the experiment, the underlying contributing factors warrant deeper consideration.
VI Conclusions